A student needed to make 150mL of a .650 M CdCl2 solution. The stock solution available had a concentration of 2.015 M. What volume of stock solution does the student need to make their solution ?
Answer #1

Using dilution effect,

M1×V1 = M2×V2

0.650×150 = 2.015×V2

V2 = 48.4L

So 48.4 mL stock solution needed